David wants throw a water ballon at george from window. George walks below window every day. With water ballon David estimates george’s Speed to be 0.5m/s. If window is 10m above ground how far in advance should David drop ballon
1
Expert's answer
2019-10-11T10:47:23-0400

The time of flight is


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(10)}{9.8}}=1.4\ s"

Thus,


"d=vt=(0.5)(1.4)=0.7\ m"